OpenStudy (anonymous):

if f(x)=2x^3+Ax^2+Bx-5 and if f(2)=3 and f(-2)=-37, what is the value of A+B? A)-6 B)-3 C)-1 D)2 E)Cant be determined

OpenStudy (anonymous):

f(2)=11 + 4 A + 2 B = 3 f(-2)=-21 + 4 A - 2 B = -37

OpenStudy (anonymous):

Solve the two equations for A and B

OpenStudy (anonymous):

You find A=-3 B=2

OpenStudy (anonymous):

A+B =-3+2=-1

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

yw